OpenStudy (anonymous):

Before a nuclear reaction, there is 0.01000 kg of uranium. After the nuclear reaction, there is 0.00950 kg of uranium. How much energy was released? (h = 6.63 × 10^-34 Js) (Points : 3) 9.00 × 10^16 J 4.50 × 10^13 J 1.50 × 10^5 J 5.00 × 10^-4 J

OpenStudy (anonymous):

the energy lost is 0.00050 by scientific notation d

OpenStudy (anonymous):

An electron is an example of which of the following? (Points : 3) antiparticle boson lepton hadron

OpenStudy (anonymous):

@leon549

OpenStudy (anonymous):

What is the energy of a single photon of red light with a wavelength of 700.0 nm? Remember that E = hf and Planck's constant is expressed by h = 6.63 × 10^-34 Js = 4.14 × 10^-15 eVs. (Points : 3) 4.29 × 10^-20 J 2.84 × 10^-19 J 4.98 × 10^-19 J 2.84 × 10^14 J

OpenStudy (anonymous):

lemda = 700nm lemda = 700 * 10 ^9 lemda = 0.700 *10 ^9

OpenStudy (anonymous):

what is lemda?

OpenStudy (anonymous):

dude can we talk it middle of the and i cant type

OpenStudy (anonymous):

its wavelenght which denoted by a simbol called lemda

OpenStudy (anonymous):

oh okay sorry

OpenStudy (anonymous):

so in order to find the frequency velocity is very imp

OpenStudy (anonymous):

How do i do it

OpenStudy (anonymous):

oh wait i Think the answer is B am I right?

OpenStudy (anonymous):

no d

OpenStudy (michele_laino):

The released energy, is given by the famous Einstein's relationship, namely: \[\Large \begin{gathered} E = \Delta m \cdot {c^2} = \left( {0.01 - 0.0095} \right) \times 9 \times {10^{16}} = \hfill \\ = 4.5 \times {10^{13}}Joules \hfill \\ \end{gathered} \]